# Leveraging Large Language Models for Career Mobility Analysis: A Study of Gender, Race, and Job Change Using U.S. Online Resume Profiles

### v1

@@ 1 Introduction
Over the past several decades, the U.S. labor market has undergone significant changes, marked by the decline of internal labor markets and the rise of external labor markets. Internal labor markets, once characterized by stable, long-term employment and opportunities for internal mobility within organizations, have diminished due to shifts in organizational structures and employment practices [14, 24]. In their place, external labor markets have expanded, driven by flexible employment arrangements and increased movement across organizations [5, 15]. As a result, employers now rely more heavily on hiring external candidates, creating multiple pathways for career mobility.
At the same time, the educational composition of the U.S. workforce has undergone dramatic changes. The labor force becoming increasingly educated over time. The share of workers with at least a college degree has risen from 26% in 1991 to nearly 42% in 2022 [47]. This shift reflects expanded access to higher education and growing demand for high-skilled workers.
@@
The college expansion has also brought about increased gender and racial/ethnic diversity in the workforce [34]. Between 1980 and 2023, the share of women and ethnoracial minority workers in high-skilled workforce has risen dramatically [42]. However, little is known about how the frequency and outcomes of internal and external job mobility vary by gender and race. On one hand, persistent structural barriers may limit the opportunities and benefits of job mobility for women and minority groups, compounding existing disadvantages. On the other hand, the increasing flexibility of the labor market may offer them new opportunities for upward mobility, particularly through external job changes that allow individuals to bypass entrenched workplace barriers and leverage external opportunities.
@@
Our study addresses the above obstacle through the use of an online resume dataset. This data source enables us to measure various types of job mobility – both internal and external – across a large sample of college graduates. Crucially, it provides detailed job histories, allowing us to construct longitudinal employment records for individual workers, unlike studies relying on LinkedIn’s aggregate audience estimates [18, 25, 2, 23]. This approach not only offers a clearer picture of job mobility patterns but also provides new insights into how these patterns interact with gender and racial/ethnic identities.
While the online resume dataset contains very detailed job mobility data, two additional challenges must be addressed: gender and race prediction and occupation label inference. As a secondary research dataset, resumes often do not contain all the essential attribute labels. Attributes such as gender and race of workers, as well as occupational labels of jobs are often missing. However, these attributes are essential for analyzing job mobility patterns.
@@
Developing robust occupation classification models is however challenging due to the scarcity of high-quality labeled data and the complexity of SOC taxonomies. To address this limitation, we present a novel few-shot [^1] Few-shot (also known as $k$-shot) prompting refers to a technique where a large language model (LLM) is provided with a small number of $k$ examples in the input prompt to guide it in performing a task. This contrasts with zero-shot prompting, where no examples are given, and the LLM relies solely on its pre-existing knowledge to perform a task. In general, few-shot prompting has been shown to significantly improve task performance compared to zero-shot prompting, as it allows the model to better generalize from limited input [7]. prompting framework for SOC classification titled FewSOC, which leverages the vast knowledge and reasoning capabilities of large language models (LLMs) like OpenAI’s GPT models. Our crowdsourced evaluation demonstrates that the SOC labels generated by FewSOC are significantly more accurate than the default labels provided with the dataset. As a result, we use the improved FewSOC-generated SOC labels in our analysis of career mobility, ensuring more reliable and valid findings.
@@ 2.1 Determinants of Career Mobility
Job mobility is a key pathway to achieving overall career advancement. However, the likelihood of job mobility is not evenly distributed across all groups. Traditionally disadvantaged populations, such as women and minorities, are less likely to experience mobility, partly due to entrenched biases that hinder their promotion opportunities (internal mobility) and compromise their future employment prospects (external mobility). These barriers constrain the career advancement of minority workers both within and across organizations [3, 35, 36, 48].
@@
Another factor contributing to these challenges is the segregation of women and minorities within restricted social networks. These networks often provide limited access to mainstream connections that are crucial for obtaining quality job information, securing referrals, and influencing hiring decisions. As a result, their opportunities for external mobility are diminished [22, 32, 38]. Additionally, women and minority workers are disproportionately concentrated in minority-dominated jobs, which typically offer lower pay and fewer opportunities for upward mobility [20].
@@
Despite well-documented barriers, little is known about how these dynamics affect highly-educated workers. While white-collar minority professionals may have better mobility prospects than their blue-collar counterparts, they often face a gendered or racialized “glass ceiling” [13]. However, the advantages of higher education may also mitigate some barriers, leading to different outcomes for this group. Even when women and minorities achieve job mobility, it is unclear if it results in upward mobility. For those starting in disadvantaged positions, mobility can provide an opportunity to move beyond the lower rungs of the occupational hierarchy, but benefits vary. Women, minorities, and lower-skilled workers face greater barriers and often experience diminished returns from job changes [8].
@@
Several factors contribute to these disparities. Racial discrimination and segregated social networks, for example, limit the gains minority workers can achieve through job mobility. In external labor markets, hiring and wage-setting discrimination is exacerbated by incomplete information about a worker’s productivity [4]. Employers may devalue the credentials of female and minority workers and attach biased social meanings to their job mobility [16]. As a result, minority workers who pursue new employment opportunities may receive lower wage offers or be funneled into lower-level positions with limited prospects for advancement. Additionally, segregated social networks further exacerbate these disadvantages by restricting access to high-quality job referrals and information [12, 39, 45]. Ultimately, job mobility for women and minorities may often result in inefficient churning between roles of similar status and pay, whether within or across organizations.
@@ 2.2 Predicting Occupations from Textual Data
Occupation is an important demographic attribute in social and computational social science research [19, 18, 25, 49, 2, 31, 17, 23]. Previous research has explored various methodologies and data sources for extracting and classifying occupations from text. Simple dictionary-based methods [49, 31, 17] were often employed due to the limited performance of traditional machine-learning classifiers [37]. For machine-learning based approaches, most studies employ traditional multi-class classification models, relying on feature engineering and machine-learning algorithms. Common feature engineering techniques include TF-IDF weighted bag-of-words model, doc2vec, and word2vec [46, 33, 6]. Supervised, unsupervised, and ensemble learning algorithms, such as SVM, logistic regression, kNN, and random forests, are widely used for classification [46, 33, 40, 21]. Additionally, more recent approaches leverage language models [29, 1]. For instance, Li et al. [29] proposed LLM4Jobs, a zero-shot query and retrieval framework utilizing Vicuna-33b. Achananuparp et. al. [1] introduced a multi-stage few-short prompting framework for occupation classification, leveraging LLMs for inferring and reranking.
@@
These studies utilize different types of data. Some researchers classify occupations using only job titles [6, 21], while others base their classification on entire job postings [29, 46, 33]. Occupation labels are typically drawn from existing standard occupational classification (SOC) taxonomies with varying levels of granularity, such as ISCO [46, 6], ESCO [29], O*NET-SOC [21], and others [41, 33, 40]. For instance, Javed et al. [21] used 23 major occupational classes from O*NET-SOC, while Li et al. [29] and Varelas et al. [46] employed up to 1,000 ESCO codes and 230 ISCO codes, respectively. Other studies, like Russ et al. [41], use 840 unique 6-digit SOC codes from the U.S. SOC 2010 system, while Mukherjee et al. (2021) focused on the five most common SOC codes in their dataset. Most datasets used in these studies are not publiclly available for research, with the exception of Li et al. [29], which released a synthetic dataset generated by LLMs.
@@
Methodologically, our FewSOC framework shares the most similarity with LLM4Jobs [29] and TGRE [1]. These methods require no training data and employ a multi-stage pipeline. In the first stage, LLMs generate candidate SOC codes for the test samples. In the second stage, a retrieval module in LLM4Jobs and TGRE, or a SOC mapping and matching module in FewSOC, identifies the most relevant SOC codes from the taxonomy. Unlike LLM4Jobs, which uses zero-shot prompting to classify one job post at a time, FewSOC employs few-shot examples to classify job title-company name pairs. Generally, few-shot prompting tends to enhance LLM performance in classification tasks compared to zero-shot prompting. Compared to TGRE, which incorporates LLM-based reranking in a third stage, FewSOC adopts a more streamlined pipeline by omitting the reranking module. In addition, FewSOC efficiently leverages batch prompting to process multiple test samples per batch, thereby reducing token usage and processing costs. This optimization enhances efficiency while introducing only a minor trade-off in classification accuracy, making FewSOC particularly suitable for large-scale datasets involving hundreds of thousands of resume profiles used in this study.
@@
To our knowledge, this work is among the few studies to utilize over 1,000 fine-grained SOC codes (at the 6-digit level or higher) from the O*NET-SOC taxonomy as class labels for multi-class classification [1]. Furthermore, we are the first to investigate the feasibility of crowdsourced SOC code annotation at scale, leveraging a large dataset of more than 10K job titles and company names sourced from online resume profiles.
@@ 3.5 Attribute Value Extraction
Gender and Race: Each profile is categorized as either male or female for gender and classified into one of the following races: White, Black, Asian, or Hispanic. To infer these attributes without identifying the workers, we utilized a separate dataset provided by Lightcast, containing only unique names extracted from the same set of profiles in our dataset. We applied our gender and race classification models to this set of names and shared the prediction results with Lightcast. They then added the inferred gender and race values to the corresponding profiles and then sent the enriched anonymized profiles (without names) back to us. Gender was inferred using a maximum likelihood model implemented with the Python package chicksexer [^5] https://github.com/kensk8er/chicksexer, achieving a testing accuracy of 93.4% [27]. Race was inferred using a logistic regression classifier trained on multiple full-name datasets, yielding a testing accuracy score of 84%.
@@
Cohort: We classify individuals into generational cohorts based on their inferred birth year, calculated by subtracting 23 (the average age of graduation) from the year they obtained their Bachelor’s degree. To determine generation, we use definitions from the Pew Research Center [10]. For instance, individuals born between 1981 and 1996 are classified as Millennials, while those born between 1965 and 1980 are categorized as Generation X. Belonging to different cohorts can be linked to career mobility in various ways, because of different economic conditions at job market entry and distinct attitudes toward career advancement.
@@ 4 Occupation Classification using Large Language Models
Fine-grained occupation classification, encompassing over a thousand unique SOC codes and titles, is limited by the scarcity of open-source labeled training data. Without sufficient labeled examples, traditional supervised learning methods often struggle to achieve the accuracy required for reliable career mobility analysis. Meanwhile, large language models (LLMs) have exhibited strong capabilities across a variety of natural language processing (NLP) tasks, due in part to their in-context learning (ICL) abilities [7]. This allows LLMs to adapt to new tasks with only a few examples, making them a promising option for addressing these data-scarce challenges.
@@ 4.1 FewSOC: Few-shot Batch Prompting Framework for SOC Classification
To minimize token usage and processing time when handling millions of job records from the Lightcast dataset, we generate responses for multiple job records per inference batch, similar to the batch prompting approach [11]. The prompt template is shown in E.
Since LLMs have limited precision in generating exact codes from large taxonomies [44, 28], the SOC mapping and matching step addresses any invalid SOC codes produced in the initial SOC generation step – those not found in the O*NET-SOC 2019 taxonomy. Invalid SOC codes may arise from SOC codes belonging to older taxonomy versions or from non-existent SOC codes. For each invalid SOC code, we iteratively look up O*NET-SOC crosswalk tables to map older SOC versions to the 2019 version. If the process results in multiple possible matches (1-to-many mappings), we then prompt the LLM to perform a batch selection of the most likely SOC code. The prompt template can be found in E. Finally, for each remaining invalid SOC code not found in the crosswalk tables, we measure its similarity with all 1,016 8-digit SOC codes in the O*NET-SOC 2019 taxonomy based on the proportion of word overlap between their titles (in Algorithm 2), replacing it with the closest match.
@@ 7 Limitations and Future Work
In-context examples: In this work, we manually engineered a fixed set of in-context examples for FewSOC. While this approach yielded substantial improvements over Lightcast’s SOC predictions, it may not guarantee optimal performance and could lead to variability in the quality and representativeness of the examples. Moreover, the manual selection method is time-consuming and inherently subjective. Future research should explore more systematic methods for selecting in-context examples, such as kNN-based approaches [30], to further enhance the effectiveness of FewSOC.
@@
Representation bias: The study utilizes online resume profiles which provide a sample of online job seekers and may not accurately represent the entire U.S. workforce. Individuals with higher education levels, strong online presences, those with well-established careers, or those actively seeking employment may be overrepresented. Online job searches have played an increasingly central role in how people find work, especially for college graduates, surpassing other means such as personal or professional connections [43]. An estimated 60% to 70% of all job vacancies are posted online, with a disproportionate number of them for high-skilled jobs [9]. Almost 80% of Americans have searched for jobs online in the last two years [43]. Therefore, we expect the online resume data to represent the highly-educated workforce reasonably well, which is our focus. Future research could consider addressing potential representation biases from online sources, particularly when studying a broader population.

### v2

@@ 1 Introduction
Over the past several decades, the U.S. labor market has undergone significant changes, marked by the decline of internal labor markets and the rise of external labor markets. Internal labor markets, once characterized by stable, long-term employment and opportunities for internal mobility within organizations, have diminished due to shifts in organizational structures and employment practices [diprete2002internal, kalleberg2000nonstandard]. In their place, external labor markets have expanded, driven by flexible employment arrangements and increased movement across organizations [bidwell2013employment, farber2008employment]. As a result, employers now rely more heavily on hiring external candidates, creating multiple pathways for career mobility.
At the same time, the educational composition of the U.S. workforce has undergone dramatic changes. The labor force has become increasingly educated over time. The share of workers with at least a college degree has risen from 26% in 1991 to nearly 42% in 2022 [brundage2017educational]. This shift reflects expanded access to higher education and growing demand for high-skilled workers.
@@
The college expansion has also brought about increased gender and racial/ethnic diversity in the workforce [ortman2009united]. Between 1980 and 2023, the share of women and ethnoracial minority workers in high-skilled workforce has risen dramatically [pew2024gendergains]. However, little is known about how the frequency and outcomes of internal and external job mobility vary by gender and race. On one hand, persistent structural barriers may limit the opportunities and benefits of job mobility for women and minority groups, compounding existing disadvantages. On the other hand, the increasing flexibility of the labor market may offer them new opportunities for upward mobility, particularly through external job changes that allow individuals to bypass entrenched workplace barriers and leverage external opportunities.
@@
Our study addresses the above obstacle through the use of an online resume dataset. This data source enables us to measure various types of job mobility – both internal and external – across a large sample of college graduates. Crucially, it provides detailed job histories, allowing us to construct longitudinal employment records for individual workers, unlike studies relying on LinkedIn’s aggregate audience estimates [haranko2018professional, kashyap2021analysing, berte2023monitoring, kalhor2024gender]. This approach not only offers a clearer picture of job mobility patterns but also provides new insights into how these patterns interact with gender and racial/ethnic identities.
While the online resume dataset contains very detailed job mobility data, two additional challenges must be addressed: gender and race prediction and occupation label inference. As a secondary research dataset, resumes often do not contain all the essential attribute labels. Attributes such as gender and race of workers, as well as occupational labels for jobs are often missing. However, these attributes are essential for analyzing job mobility patterns.
@@
Developing robust occupation classification models is however challenging due to the scarcity of high-quality labeled data and the complexity of SOC taxonomies. To address this limitation, we present a novel few-shot [^1] Few-shot (also known as $k$-shot) prompting refers to a technique where a large language model (LLM) is provided with a small number of $k$ examples in the input prompt to guide it in performing a task. This contrasts with zero-shot prompting, where no examples are given, and the LLM relies solely on its pre-existing knowledge to perform a task. In general, few-shot prompting has been shown to significantly improve task performance compared to zero-shot prompting, as it allows the model to better generalize from limited input [brown2020language]. prompting framework for SOC classification titled FewSOC, which leverages the vast knowledge and reasoning capabilities of large language models (LLMs) like OpenAI’s GPT models. Our crowdsourced evaluation demonstrates that the SOC labels generated by FewSOC are significantly more accurate than the default labels provided with the dataset. As a result, we use the improved FewSOC-generated SOC labels in our analysis of career mobility, ensuring more reliable and valid findings.
@@ 2.1 Determinants of Career Mobility
Job mobility is a key pathway to achieving overall career advancement. However, the likelihood of job mobility is not evenly distributed across all groups. Traditionally disadvantaged populations, such as women and minorities, are less likely to experience mobility, partly due to entrenched biases that hinder their promotion opportunities (internal mobility) and compromise their future employment prospects (external mobility). These barriers constrain the career advancement of minority workers both within and across organizations [bertrand2004emily, pager2003mark, pager2009discrimination, waldinger2003other].
@@
Another factor contributing to these challenges is the segregation of women and minorities within restricted social networks. These networks often provide limited access to mainstream connections that are crucial for obtaining quality job information, securing referrals, and influencing hiring decisions. As a result, their opportunities for external mobility are diminished [kalfa2018social, moss2001stories, petersen2000offering]. Additionally, women and minority workers are disproportionately concentrated in minority-dominated jobs, which typically offer lower pay and fewer opportunities for upward mobility [huffman2004racial].
@@
Despite well-documented barriers, little is known about how these dynamics affect highly-educated workers. While white-collar minority professionals may have better mobility prospects than their blue-collar counterparts, they often face a gendered or racialized “glass ceiling” [cox1990invisible]. However, the advantages of higher education may also mitigate some barriers, leading to different outcomes for this group. Even when women and minorities achieve job mobility, it is unclear if it results in upward mobility. For those starting in disadvantaged positions, mobility can provide an opportunity to move beyond the lower rungs of the occupational hierarchy, but benefits vary. Women, minorities, and lower-skilled workers face greater barriers and often experience diminished returns from job changes [bukodi2010bad].
@@
Several factors contribute to these disparities. Racial discrimination and segregated social networks, for example, limit the gains minority workers can achieve through job mobility. In external labor markets, hiring and wage-setting discrimination is exacerbated by incomplete information about a worker’s productivity [bidwell2011paying]. Employers may devalue the credentials of female and minority workers and attach biased social meanings to their job mobility [fuller2008job]. As a result, minority workers who pursue new employment opportunities may receive lower wage offers or be funneled into lower-level positions with limited prospects for advancement. Additionally, segregated social networks further exacerbate these disadvantages by restricting access to high-quality job referrals and information [collins1997black, royster2003race, stainback2008social]. Ultimately, job mobility for women and minorities may often result in inefficient churning between roles of similar status and pay, whether within or across organizations.
@@ 2.2 Predicting Occupations from Textual Data
Occupation is an important demographic attribute in social and computational social science research [hu2016language, haranko2018professional, kashyap2021analysing, zhao2021exploring, berte2023monitoring, miyazaki2024public, guo2024evolution, kalhor2024gender]. Previous research has explored various methodologies and data sources for extracting and classifying occupations from text. Simple dictionary-based methods [zhao2021exploring, miyazaki2024public, guo2024evolution] were often employed due to the limited performance of traditional machine-learning classifiers [pan2019twitter]. For machine-learning based approaches, most studies employ traditional multi-class classification models, relying on feature engineering and machine-learning algorithms. Common feature engineering techniques include TF-IDF weighted bag-of-words model, doc2vec, and word2vec [varelas2022employing, mukherjee2021determining, boselli2017using]. Supervised, unsupervised, and ensemble learning algorithms, such as SVM, logistic regression, kNN, and random forests, are widely used for classification [varelas2022employing, mukherjee2021determining, russ2016computer, javed2015carotene]. Additionally, more recent approaches leverage language models [li2023llm4jobs, achananuparp2025multi]. For instance, Li et al. [li2023llm4jobs] proposed LLM4Jobs, a zero-shot query and retrieval framework utilizing Vicuna-33b. Achananuparp et al. [achananuparp2025multi] introduced a multi-stage few-short prompting framework for occupation classification, leveraging LLMs for inferring and reranking.
@@
These studies utilize different types of data. Some researchers classify occupations using only job titles [boselli2017using, javed2015carotene], while others base their classification on entire job postings [li2023llm4jobs, varelas2022employing, mukherjee2021determining]. Occupation labels are typically drawn from existing standard occupational classification (SOC) taxonomies with varying levels of granularity, such as ISCO [varelas2022employing, boselli2017using], ESCO [li2023llm4jobs], O*NET-SOC [javed2015carotene], and others [russ2023evaluation, mukherjee2021determining, russ2016computer]. For instance, Javed et al. [javed2015carotene] used 23 major occupational classes from O*NET-SOC, while Li et al. [li2023llm4jobs] and Varelas et al. [varelas2022employing] employed up to 1,000 ESCO codes and 230 ISCO codes, respectively. Other studies, like Russ et al. [russ2023evaluation], use 840 unique 6-digit SOC codes from the U.S. SOC 2010 system, while Mukherjee et al. (2021) focused on the five most common SOC codes in their dataset. Most datasets used in these studies are not publicly available for research, with the exception of Li et al. [li2023llm4jobs], which released a synthetic dataset generated by LLMs.
@@
Methodologically, our FewSOC framework shares the most similarity with LLM4Jobs [li2023llm4jobs] and TGRE [achananuparp2025multi]. These methods require no training data and employ a multi-stage pipeline. In the first stage, LLMs generate candidate SOC codes for the test samples. In the second stage, a retrieval module in LLM4Jobs and TGRE, or a SOC mapping and matching module in FewSOC, identifies the most relevant SOC codes from the taxonomy. Unlike LLM4Jobs, which uses zero-shot prompting to classify one job post at a time, FewSOC employs few-shot examples to classify job title-company name pairs. Generally, few-shot prompting tends to enhance LLM performance in classification tasks compared to zero-shot prompting. Compared to TGRE, which incorporates LLM-based reranking in a third stage, FewSOC adopts a more streamlined pipeline by omitting the reranking module. In addition, FewSOC efficiently leverages batch prompting to process multiple test samples per batch, thereby reducing token usage and processing costs. This optimization enhances efficiency while introducing only a minor trade-off in classification accuracy, making FewSOC particularly suitable for large-scale datasets involving hundreds of thousands of resume profiles used in this study.
@@
To our knowledge, this work is among the few studies to utilize over 1,000 fine-grained SOC codes (at the 6-digit level or higher) from the O*NET-SOC taxonomy as class labels for multi-class classification [achananuparp2025multi]. Furthermore, we are the first to investigate the feasibility of crowdsourced SOC code annotation at scale, leveraging a large dataset of more than 10K job titles and company names sourced from online resume profiles.
@@ 3.5 Attribute Value Extraction
Gender and Race: Each profile is categorized as either male or female for gender and classified into one of the following races: White, Black, Asian, or Hispanic. To infer these attributes without identifying the workers, we utilized a separate dataset provided by Lightcast, containing only unique names extracted from the same set of profiles in our dataset. We applied our gender and race classification models to this set of names and shared the prediction results with Lightcast. They then added the inferred gender and race values to the corresponding profiles and then sent the enriched anonymized profiles (without names) back to us. Gender was inferred using a maximum likelihood model implemented with the Python package chicksexer [^5] https://github.com/kensk8er/chicksexer, achieving a testing accuracy of 93.4% [krstovski2023computation]. Race was inferred using a logistic regression classifier trained on multiple full-name datasets, yielding a testing accuracy score of 84%.
@@
Cohort: We classify individuals into generational cohorts based on their inferred birth year, calculated by subtracting 23 (the average age of graduation) from the year they obtained their Bachelor’s degree. To determine generation, we use definitions from the Pew Research Center [pew2015]. For instance, individuals born between 1981 and 1996 are classified as Millennials, while those born between 1965 and 1980 are categorized as Generation X. Belonging to different cohorts can be linked to career mobility in various ways, because of different economic conditions at job market entry and distinct attitudes toward career advancement.
@@ 4 Occupation Classification using Large Language Models
Fine-grained occupation classification, encompassing over a thousand unique SOC codes and titles, is limited by the scarcity of open-source labeled training data. Without sufficient labeled examples, traditional supervised learning methods often struggle to achieve the accuracy required for reliable career mobility analysis. Meanwhile, large language models (LLMs) have exhibited strong capabilities across a variety of natural language processing (NLP) tasks, due in part to their in-context learning (ICL) abilities [brown2020language]. This allows LLMs to adapt to new tasks with only a few examples, making them a promising option for addressing these data-scarce challenges.
@@ 4.1 FewSOC: Few-shot Batch Prompting Framework for SOC Classification
To minimize token usage and processing time when handling millions of job records from the Lightcast dataset, we generate responses for multiple job records per inference batch, similar to the batch prompting approach [cheng2023batch]. The prompt template is shown in E.
Since LLMs have limited precision in generating exact codes from large taxonomies [soroush2024large, lee2024large], the SOC mapping and matching step addresses any invalid SOC codes produced in the initial SOC generation step – those not found in the O*NET-SOC 2019 taxonomy. Invalid SOC codes may arise from SOC codes belonging to older taxonomy versions or from non-existent SOC codes. For each invalid SOC code, we iteratively look up O*NET-SOC crosswalk tables to map older SOC versions to the 2019 version. If the process results in multiple possible matches (1-to-many mappings), we then prompt the LLM to perform a batch selection of the most likely SOC code. The prompt template can be found in E. Finally, for each remaining invalid SOC code not found in the crosswalk tables, we measure its similarity to all 1,016 8-digit SOC codes in the O*NET-SOC 2019 taxonomy based on the proportion of word overlap between their titles (in Algorithm 2), replacing it with the closest match.
@@ 7 Limitations and Future Work
In-context examples: In this work, we manually engineered a fixed set of in-context examples for FewSOC. While this approach yielded substantial improvements over Lightcast’s SOC predictions, it may not guarantee optimal performance and could lead to variability in the quality and representativeness of the examples. Moreover, the manual selection method is time-consuming and inherently subjective. Future research should explore more systematic methods for selecting in-context examples, such as kNN-based approaches [liu2022makes], to further enhance the effectiveness of FewSOC.
@@
Representation bias: The study utilizes online resume profiles which provide a sample of online job seekers and may not accurately represent the entire U.S. workforce. Individuals with higher education levels, strong online presences, those with well-established careers, or those actively seeking employment may be overrepresented. Online job searches have played an increasingly central role in how people find work, especially for college graduates, surpassing other means such as personal or professional connections [smith2015]. An estimated 60% to 70% of all job vacancies are posted online, with a disproportionate number of them for high-skilled jobs [carnevale2014understanding]. Almost 80% of Americans have searched for jobs online in the last two years [smith2015]. Therefore, we expect the online resume data to represent the highly-educated workforce reasonably well, which is our focus. Future research could consider addressing potential representation biases from online sources, particularly when studying a broader population.